# Waterloo, Ontario, N2L 2Y5, Canada

### v1

@@ 3 The generalized $u<d$ converse
In the previous section, we have shown that the 2-to-2 CWT has a converse in pure AdS3. In this section, we demonstrate the consequences of this result for a CWT converse in general AdS3 vacuum solutions, yielding an extension of the $u<d$ result of Caminiti:2024ctd to a broader class of backgrounds.
In three spacetime dimensions, vacuum solutions of Einstein’s equations with negative cosmological constant, such as conical defect and BTZ black hole solutions, are locally AdS3 Brown:1988am , but may have nontrivial topology. Consider one such vacuum solution, $\mathcal{M}$, and consider an open, simply connected region $\mathcal{N}\subset\mathcal{M}$ which does not contain defects or singularities. Because the manifold is locally AdS3, $\mathcal{N}$ can be endowed with coordinates under which the line element is that of AdS3 in e.g., standard global coordinates; in other words, $\mathcal{N}$ is a portion of pure AdS3.
@@
    the boundary state is pure. [^7] Note that we allow for the possibility that $\mathcal{M}$ contains multiple asymptotically AdS3 boundaries. An example of holographic scattering in the multi-boundary case is the two-sided BTZ setup in Caminiti:2024ctd .
@@
The above assumptions come with two consequences which will be important for our analysis. Firstly, assumption 1 implies $\mathcal{N}$ contains a disconnected entanglement wedge candidate for $\hat{\mathcal{V}}_{1}\cup\hat{\mathcal{V}}_{2}$, with corresponding RT surface denoted $d$. Here, $u$ and $d$ are uniquely defined because in pure AdS3, each spacelike boundary interval has a unique homologous extremal surface (and $\mathcal{N}$ is an open portion of pure AdS3). In particular, the $d$ entanglement wedge is simply the union of the $\hat{\mathcal{V}}_{1}$ causal wedge and the $\hat{\mathcal{V}}_{2}$ causal wedge. Secondly, the assumption that the boundary state is pure implies $\gamma_{{\hat{\mathcal{R}}}_{1}}=\gamma_{{\hat{\mathcal{R}}}_{2}}$ by virtue of ${\hat{\mathcal{R}}}_{1}$ and ${\hat{\mathcal{R}}}_{2}$ being causal complements Hubeny:2012wa . We remind the reader that purity is not intrinsic to $\mathcal{N}$; rather, purity requires constraints on the global nature of the bulk geometry, such as the existence or nonexistence of horizons and singularities.
@@
This result generalizes and clarifies the conical defect and BTZ examples considered in Caminiti:2024ctd . For instance, Caminiti:2024ctd found that $u<d$ implies scattering in heavy conical defect geometries when ${\hat{\mathcal{V}}}_{1}$ and ${\hat{\mathcal{V}}}_{2}$ lie in a constant time-slice and are taken to be of equal size (e.g., see figure 2 below). Our result reveals $u<d$ implies scattering for any choice of spacelike separated ${\hat{\mathcal{V}}}_{1}$ and ${\hat{\mathcal{V}}}_{2}$, as a simple consequence of the local AdS3 structure of the geometry.
We emphasize that the general arguments in this section apply well beyond the static BTZ and conical defect solutions studied in Caminiti:2024ctd ; other nontrivial vacuum solutions where our results would apply include, e.g., spinning defects (e.g., Miskovic:2009uz ; Compere:2018aar ; Edelstein:2011vu ; Briceno:2021dpi ; Li:2024rma ), spinning BTZ black holes Banados:1992wn ; Banados:1992gq , and orbiting defects and black holes Balasubramanian:1999zv . Again, in these geometries, the scattering setup extends beyond that considered in Caminiti:2024ctd ; we can choose the input regions, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, to be any spacelike separated causal diamonds on the boundary cylinder while the output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ would be the causal diamonds extending between the future tips of these input regions. In appendix B, we examine the example of spinning defects in detail.
What are the limitations of this result? Recall that from assumptions 1-3, the scattering region described by the final expression in eq. (25) is necessarily contained in the $u$ entanglement wedge, and hence also in $\mathcal{N}$. As a consequence, the scattering region does not contain matter sources (which deform the local geometry). [^9] Interestingly, since all of our three assumptions are always satisfied in the heavy defect geometry (see Caminiti:2024ctd ), we can conclude that scattering regions in the defect geometry never contain the defect itself. Physically, this arises because probe systems approaching the defect at $r=0$ experience significant time delays, and large time delays enable boundary scattering to occur. This is a clue that the presence of matter sources will disrupt the relationship between extremal surfaces and bulk scattering; we will elaborate on this theme in the following section.
@@ 4.1 Shell geometry
One may ask why the resulting geometry is static if it contains a shell of matter which seemingly would want to collapse in on itself. Indeed, in appendix C, we show that the shell has positive tension which makes it want to contract. The resolution lies in recognizing that the shell functions as a domain wall separating a bubble of true vacuum with M=1M=-1 from an exterior region with higher energy, M>1M>-1. Since the interior has lower energy, the bubble wishes to expand as in Coleman:1980aw . Ultimately, the geometry remains static because the positive tension of the shell precisely balances the outward pressure exerted by the bubble.
@@ 4.3 A simple counterexample to $u<d$
To see this, imagine fixing $\mu<1/2$, [^15] From Caminiti:2024ctd , this condition ensures that without the shell, assumptions 1-3 of section 3 are obeyed. and slowly increasing the size $R$ of the shell from $R=0$. When $R$ is small, our three assumptions in section 3 are obeyed, and therefore section 3 tells us the scattering region lies somewhere inside the $u$ entanglement wedge. Now, as the shell increases in size, it may disrupt the $u$ surface without affecting the scattering region. In particular, letting $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ be associated with intervals of equal size $x$ and separation $\theta$ on a constant-time slice of the cylinder (see figure 2), we may end up in a scenario where scattering is at the conical defect threshold Caminiti:2024ctd , i.e.,
@@
To close with a brief aside, we comment that eq. (44) is nevertheless a nontrivial result. [^18] We thank the authors of ap ; ap2 for bringing this point to our attention. It is nontrivial because sometimes $o$ is the minimal candidate, and so the CWT need only imply $o<d$ (not $i<d$). If we increase $\theta+x$, we can arrive at a situation where neither $i$ nor $u$ exists, in which case the reasoning in footnote 17 ensures
@@
Inspired by ap ; ap2 , one can summarize these results by writing $u_{simple}<d$, where we define $u_{simple}$ as the unique candidate in $\{o,u,i\}$ which is contained in the simple wedge Engelhardt:2014gca ; Engelhardt:2021mue of the boundary interval with opening angle $\theta+x$. Notably, while until now we have focused on the case $\theta<\Delta\phi^{*}$, we analyze phase diagrams and holographic scattering in the shell geometry in appendix C, and we find that for $\theta>\Delta\phi^{*}$, the $u_{simple}<d$ result appears to persist on a case-by-case basis. See further discussions on this point in section 6.
@@ 5.1 The $n$-to-$n$ CWT
Let us begin by stating the $n$-to-$n$ connected wedge theorem (adapted from May:2022clu ).
@@
We comment that the ${\hat{\mathcal{R}}}_{i}$ output regions shown in figure 9 are larger than the ${\hat{\mathcal{W}}}_{i}$ output regions defined in eq. (46). One can check by the equivalence of causal and entanglement wedges in pure AdS3 that $J^{-}({\hat{\mathcal{W}}}_{i})=J^{-}(r_{i})=J^{-}({\hat{\mathcal{R}}}_{i})$, and hence we are free to perform the replacement ${\hat{\mathcal{W}}}_{i}\to{\hat{\mathcal{R}}}_{i}$ without modifying the scattering region of interest. At a deeper level, we expect that it should be possible to upgrade the $n$-to-$n$ CWT as originally stated to accommodate larger output regions, since e.g., upgrading ${\hat{\mathcal{W}}}_{i}\to{\hat{\mathcal{R}}}_{i}$ does not affect the focusing or homology arguments in the CWT proof. [^19] We thank Alex May for discussion on this topic. We emphasize that the 2-to-2 CWT of May:2021nrl (our theorem 1), accommodates output regions larger than the $\hat{\mathcal{W}}_{i}$; we used this already in figure 4.
@@ 5.4 $n$-to-$n$ in conical defect background
The case $n=2$ was considered in Caminiti:2024ctd , where a significant difference was found between the points-based and regions-based formulations of the CWT. However for our $n$-to-$n$ setup with $n\geq 3$, our input and output regions always have a width which is less than $\pi$. This observation implies that the entanglement wedge of each region equals its causal wedge Caminiti:2024ctd , and therefore the regions-based CWT is equivalent to the points-based CWT for $n\geq 3$.
@@
holographic scattering is impossible for all choices of $\theta$ and $x$. Hence, no competition among RT surfaces will suffice to imply holographic scattering. [^29] For the sake of completeness, we here include the phase transition from the fully disconnected RT candidate to the fully connected RT candidate, with $n=3$: $\sin^{2}\left(\sqrt{|M|}\frac{\theta-x}{2}\right)\sin(\sqrt{|M|}\frac{2\theta+x}{2})<\sin^{3}\left(\sqrt{|M|}\frac{x}{2}\right)\,.$ (80) This generalizes the case of 2-to-2 points-based scattering in the defect geometry, where as discussed in Caminiti:2024ctd , scattering is impossible with M>4/9M>-4/9.
@@
which is never satisfied since we always have $\theta>x$ and $x<2\pi/n$. Further, in the $\theta\leq 2\pi/n$ regime, eq. (83) is necessary for scattering since scattering to a number of output points at $t=\pi-\theta$ and one output point with $t_{r_{m}}<\pi-\theta$ is more difficult than, and hence implies, scattering to output points all at or above $t=\pi-\theta$. We conclude that in the mass regime (79), it is impossible to have holographic scattering in our $n$-to-$n$ setup, generalizing the points-based scattering result of Caminiti:2024ctd .
@@ 6 Discussion
In this paper, we explored the generality of the relationship between holographic scattering and nonminimal surfaces identified in Caminiti:2024ctd . Starting with the observation that the 2-to-2 connected wedge theorem (CWT) admits a converse in pure AdS3, we established a more general connection between holographic scattering and nonminimal surfaces in pure states for any spacetime where the geometry near the boundary is locally AdS3. This extends the $u<d$ result of Caminiti:2024ctd .
We emphasize that our general arguments apply well beyond the static BTZ and conical defect solutions considered in Caminiti:2024ctd . Other nontrivial solutions where our results would apply include, e.g., spinning defects Miskovic:2009uz ; Compere:2018aar ; Edelstein:2011vu ; Briceno:2021dpi ; Li:2024rma (see appendix B), spinning BTZ black holes Banados:1992wn ; Banados:1992gq , and orbiting or colliding defects and black holes Balasubramanian:1999zv . Further, in these spacetimes, our scattering setup can be generalized beyond that considered in Caminiti:2024ctd and illustrated in figure 2, in which $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ have equal size and lie on the same time-slice. Our general arguments allow us to choose the input regions $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ to be any spacelike separated causal diamonds on the boundary cylinder, where the output regions $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ are the causal diamonds extending between the future tips of these inputs. Consequently, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ need not be the same size or lie on a constant time slice, e.g., see figure 4.
@@
Maximal outputs: A key feature of the $u<d$ construction is that the output regions $\hat{\mathcal{R}}_{i}$ are chosen to be maximal, given fixed inputs $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$. We emphasize that if the outputs are smaller, the $u<d$ relation need not imply scattering. In particular, consider sliding the endpoints of $\hat{\mathcal{R}}_{i}$ up along the $\hat{\mathcal{J}}^{-}(r_{i})$ lightcones to produce smaller $\hat{\mathcal{W}}_{i}\subset\hat{\mathcal{R}}_{i}$ – for reference, see figure 4. It turns out that shrinking $\hat{\mathcal{W}}_{i}$ too far may result in an empty scattering region. A dramatic illustration of this situation is given by considering $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ to be of equal size, on the boundary of a conical defect geometry with M>4/9M>-4/9. If we set $\hat{\mathcal{W}}_{1}={r_{1}}$ and $\hat{\mathcal{W}}_{2}={r_{2}}$, then holographic scattering is impossible for all choices of $\theta$ and $x$, even when we have $u<d$ Caminiti:2024ctd .
@@
Closed timelike curves: Another notable feature of the $u<d$ result is that in the spinning defect backgrounds, the $u$ surface may probe a region of the geometry with closed timelike curves, even if the scattering itself happens outside of the CTC region. [^31] In the notation of appendix B, this arises when $\theta+x\in(\Delta\phi_{CTC},\pi/b_{+})$ and $\theta<\Delta\phi_{CTC}$. In physical scenarios, one expects that the spinning defect solution describes the spacetime only outside of a localized matter source, such as the spinning string of Maxfield:2022rry , and no CTC region actually exists. Hence, in these physical spacetimes, we expect that the RT surfaces will encounter the matter sources in certain regimes, disrupting the $u<d$ connection to holographic scattering, similarly to the situation in section 4. Still it would be interesting to understand, from a bulk reconstruction point of view, what it means for the $u$ entanglement wedge to probe the CTC region, and why this appears to be useful for diagnosing holographic scattering.
@@
Specifically, in geometries which are locally AdS3 outside of certain matter sources, the $u$ surface reaches deeper and deeper into the bulk as $\theta+x$ increases, and it may become disrupted by the presence of bulk matter at a critical value of $\theta+x$ which we call $\Delta\phi_{s}$. For static defect geometries with M<1/4M<-1/4, Caminiti:2024ctd found $\Delta\phi_{s}=\frac{\pi}{\sqrt{|M|}}$, and for spinning defect geometries with $M<-\frac{1}{4}-J^{2}$, we generalized this to
@@
Our goal in studying the shell geometry in section 4 was to further examine the effects of matter on the $u<d$ converse. In this case, we have $\Delta\phi_{s}=\frac{2}{\mu}\tan^{-1}\frac{\mu}{R}$, and we found examples where holographic scattering is controlled by a naive extension of the $u=d$ formula and yet $\theta+x>\Delta\phi_{s}$. Hence in these situations, the $u$ surface does not exist and so cannot be said to control the scattering. This reinforces the result that bulk matter disrupts the tight relationship between extremal surfaces and bulk scattering introduced in Caminiti:2024ctd .
@@
Conversely, one might conjecture that whenever $u<d$ fails to control scattering, the $u$ surface does not exist at the scattering threshold. This is almost implied by our discussion of eq. (85), except for the possibility that the $\theta+x=\Delta\phi_{s}$ curve intersects the scattering curve not only at eq. (85) but also at some $\theta$ above $\Delta\phi^{*}$, which is the regime where scattering is no longer controlled by the naive extension of the $u=d$ formula (see figure 16). To validate this conjecture in cases where $R$ is very large ($R>>1$), or the shell is very light ($\mu\approx 1$), note that $\Delta\phi_{s}\approx\Delta\phi^{*}$ in these limits. [^33] To be more quantitative: for $R>1$, we find $\Delta\phi_{s}-\Delta\phi^{*}<\pi/20\simeq 0.157$ for all choices of $\mu$. The same bound holds with $\mu>0.9$ and any choice of $R$. Hence, for $\theta>\Delta\phi^{*}$, the $u$ surface only exists ($\theta<\Delta\phi_{s}-x$) within a tiny island of parameter space near $x=0$, and it is reasonable to expect that this island does not extend to the scattering threshold curve. Turning now to cases where $R$ is small and the shell is heavy, we observe that in these cases, the (extended) $u=d$ curve provides a lower bound on the scattering threshold, since the shell introduces time delays which make scattering more difficult compared to scattering in the $R=0$ geometry. [^34] Interestingly, this feature does not hold for all parameters; with the introduction of a sufficiently large shell, it is possible that scattering becomes easier – see figure 16. This is because, roughly, light rays travel faster through pure AdS3 than they do in the conical defect geometry, and so a large pure AdS3 region can compensate for the time delays introduced at the shell interface. In these cases, the conjecture is validated by eq. (85) being less than $\Delta\phi^{*}$. To fully validate this conjecture, one must be more careful about intermediate values of $R$, $\mu$, but it appears to hold on a case-by-case basis (e.g., $R=0.75$, $\mu=0.4$).

### v3

@@ 3 The generalized $u<d$ converse
In the previous section, we have shown that the 2-to-2 CWT has a converse in pure AdS3. In this section, we demonstrate the consequences of this result for a CWT converse in general AdS3 vacuum solutions, yielding an extension of the $u<d$ result of Caminiti et al. (2024) to a broader class of backgrounds.
In three spacetime dimensions, vacuum solutions of Einstein’s equations with negative cosmological constant, such as conical defect and BTZ black hole solutions, are locally AdS3 Brown (1988), but may have nontrivial topology. Consider one such vacuum solution, $\mathcal{M}$, and consider an open, simply connected region $\mathcal{N}\subset\mathcal{M}$ which does not contain defects or singularities. Because the manifold is locally AdS3, $\mathcal{N}$ can be endowed with coordinates under which the line element is that of AdS3 in e.g., standard global coordinates; in other words, $\mathcal{N}$ is a portion of pure AdS3.
@@
    the boundary state is pure. [^7] Note that we allow for the possibility that $\mathcal{M}$ contains multiple asymptotically AdS3 boundaries. An example of holographic scattering in the multi-boundary case is the two-sided BTZ setup in Caminiti et al. (2024).
@@
The above assumptions come with two consequences which will be important for our analysis. Firstly, assumption 1 implies $\mathcal{N}$ contains a disconnected entanglement wedge candidate for $\hat{\mathcal{V}}_{1}\cup\hat{\mathcal{V}}_{2}$, with corresponding RT surface denoted $d$. Here, $u$ and $d$ are uniquely defined because in pure AdS3, each spacelike boundary interval has a unique homologous extremal surface (and $\mathcal{N}$ is an open portion of pure AdS3). In particular, the $d$ entanglement wedge is simply the union of the $\hat{\mathcal{V}}_{1}$ causal wedge and the $\hat{\mathcal{V}}_{2}$ causal wedge. Secondly, the assumption that the boundary state is pure implies $\gamma_{{\hat{\mathcal{R}}}_{1}}=\gamma_{{\hat{\mathcal{R}}}_{2}}$ by virtue of ${\hat{\mathcal{R}}}_{1}$ and ${\hat{\mathcal{R}}}_{2}$ being causal complements Hubeny and Rangamani (2012). We remind the reader that purity is not intrinsic to $\mathcal{N}$; rather, purity requires constraints on the global nature of the bulk geometry, such as the existence or nonexistence of horizons and singularities.
@@
This result generalizes and clarifies the conical defect and BTZ examples considered in Caminiti et al. (2024). For instance, Caminiti et al. (2024) found that $u<d$ implies scattering in heavy conical defect geometries when ${\hat{\mathcal{V}}}_{1}$ and ${\hat{\mathcal{V}}}_{2}$ lie in a constant time-slice and are taken to be of equal size (e.g., see figure 2 below). Our result reveals $u<d$ implies scattering for any choice of spacelike separated ${\hat{\mathcal{V}}}_{1}$ and ${\hat{\mathcal{V}}}_{2}$, as a simple consequence of the local AdS3 structure of the geometry.
We emphasize that the general arguments in this section apply well beyond the static BTZ and conical defect solutions studied in Caminiti et al. (2024); other nontrivial vacuum solutions where our results would apply include, e.g., spinning defects (e.g., Miskovic and Zanelli (2009); Compère and Fiorucci (2018); Edelstein et al. (2011); Briceño et al. (2021); Li (2024)), spinning BTZ black holes Banados et al. (1992, 1993), and orbiting defects and black holes Balasubramanian and Ross (2000). Again, in these geometries, the scattering setup extends beyond that considered in Caminiti et al. (2024); we can choose the input regions, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$, to be any spacelike separated causal diamonds on the boundary cylinder while the output regions, $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ would be the causal diamonds extending between the future tips of these input regions. In appendix B, we examine the example of spinning defects in detail.
What are the limitations of this result? Recall that from assumptions 1-3, the scattering region described by the final expression in eq. (25) is necessarily contained in the $u$ entanglement wedge, and hence also in $\mathcal{N}$. As a consequence, the scattering region does not contain matter sources (which deform the local geometry). [^9] Interestingly, since all of our three assumptions are always satisfied in the heavy defect geometry (see Caminiti et al. (2024)), we can conclude that scattering regions in the defect geometry never contain the defect itself. Physically, this arises because probe systems approaching the defect at $r=0$ experience significant time delays, and large time delays enable boundary scattering to occur. This is a clue that the presence of matter sources will disrupt the relationship between extremal surfaces and bulk scattering; we will elaborate on this theme in the following section.
@@ 4.1 Shell geometry
One may ask why the resulting geometry is static if it contains a shell of matter which seemingly would want to collapse in on itself. Indeed, in appendix C, we show that the shell has positive tension which makes it want to contract. The resolution lies in recognizing that the shell functions as a domain wall separating a bubble of true vacuum with M=1M=-1 from an exterior region with higher energy, M>1M>-1. Since the interior has lower energy, the bubble wishes to expand as in Coleman and De Luccia (1980). Ultimately, the geometry remains static because the positive tension of the shell precisely balances the outward pressure exerted by the bubble.
@@ 4.3 A simple counterexample to $u<d$
To see this, imagine fixing $\mu<1/2$, [^15] From Caminiti et al. (2024), this condition ensures that without the shell, assumptions 1-3 of section 3 are obeyed. and slowly increasing the size $R$ of the shell from $R=0$. When $R$ is small, our three assumptions in section 3 are obeyed, and therefore section 3 tells us the scattering region lies somewhere inside the $u$ entanglement wedge. Now, as the shell increases in size, it may disrupt the $u$ surface without affecting the scattering region. In particular, letting $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ be associated with intervals of equal size $x$ and separation $\theta$ on a constant-time slice of the cylinder (see figure 2), we may end up in a scenario where scattering is at the conical defect threshold Caminiti et al. (2024), i.e.,
@@
To close with a brief aside, we comment that eq. (44) is nevertheless a nontrivial result. [^18] We thank the authors of Lima et al. (2025b, a) for bringing this point to our attention. It is nontrivial because sometimes $o$ is the minimal candidate, and so the CWT need only imply $o<d$ (not $i<d$). If we increase $\theta+x$, we can arrive at a situation where neither $i$ nor $u$ exists, in which case the reasoning in footnote 17 ensures
@@
Inspired by Lima et al. (2025b, a), one can summarize these results by writing $u_{simple}<d$, where we define $u_{simple}$ as the unique candidate in $\{o,u,i\}$ which is contained in the simple wedge Engelhardt and Wall (2015); Engelhardt et al. (2021) of the boundary interval with opening angle $\theta+x$. Notably, while until now we have focused on the case $\theta<\Delta\phi^{*}$, we analyze phase diagrams and holographic scattering in the shell geometry in appendix C, and we find that for $\theta>\Delta\phi^{*}$, the $u_{simple}<d$ result appears to persist on a case-by-case basis. See further discussions on this point in section 6.
@@ 5.1 The $n$-to-$n$ CWT
Let us begin by stating the $n$-to-$n$ connected wedge theorem (adapted from May et al. (2022)).
@@
We comment that the ${\hat{\mathcal{R}}}_{i}$ output regions shown in figure 9 are larger than the ${\hat{\mathcal{W}}}_{i}$ output regions defined in eq. (46). One can check by the equivalence of causal and entanglement wedges in pure AdS3 that $J^{-}({\hat{\mathcal{W}}}_{i})=J^{-}(r_{i})=J^{-}({\hat{\mathcal{R}}}_{i})$, and hence we are free to perform the replacement ${\hat{\mathcal{W}}}_{i}\to{\hat{\mathcal{R}}}_{i}$ without modifying the scattering region of interest. At a deeper level, we expect that it should be possible to upgrade the $n$-to-$n$ CWT as originally stated to accommodate larger output regions, since e.g., upgrading ${\hat{\mathcal{W}}}_{i}\to{\hat{\mathcal{R}}}_{i}$ does not affect the focusing or homology arguments in the CWT proof. [^19] We thank Alex May for discussion on this topic. We emphasize that the 2-to-2 CWT of May (2021) (our theorem 1), accommodates output regions larger than the $\hat{\mathcal{W}}_{i}$; we used this already in figure 4.
@@ 5.4 $n$-to-$n$ in conical defect background
The case $n=2$ was considered in Caminiti et al. (2024), where a significant difference was found between the points-based and regions-based formulations of the CWT. However for our $n$-to-$n$ setup with $n\geq 3$, our input and output regions always have a width which is less than $\pi$. This observation implies that the entanglement wedge of each region equals its causal wedge Caminiti et al. (2024), and therefore the regions-based CWT is equivalent to the points-based CWT for $n\geq 3$.
@@
holographic scattering is impossible for all choices of $\theta$ and $x$. Hence, no competition among RT surfaces will suffice to imply holographic scattering. [^29] For the sake of completeness, we here include the phase transition from the fully disconnected RT candidate to the fully connected RT candidate, with $n=3$: $\sin^{2}\left(\sqrt{|M|}\frac{\theta-x}{2}\right)\sin(\sqrt{|M|}\frac{2\theta+x}{2})<\sin^{3}\left(\sqrt{|M|}\frac{x}{2}\right)\,.$ (80) This generalizes the case of 2-to-2 points-based scattering in the defect geometry, where as discussed in Caminiti et al. (2024), scattering is impossible with M>4/9M>-4/9.
@@
which is never satisfied since we always have $\theta>x$ and $x<2\pi/n$. Further, in the $\theta\leq 2\pi/n$ regime, eq. (83) is necessary for scattering since scattering to a number of output points at $t=\pi-\theta$ and one output point with $t_{r_{m}}<\pi-\theta$ is more difficult than, and hence implies, scattering to output points all at or above $t=\pi-\theta$. We conclude that in the mass regime (79), it is impossible to have holographic scattering in our $n$-to-$n$ setup, generalizing the points-based scattering result of Caminiti et al. (2024).
@@ 6 Discussion
In this paper, we explored the generality of the relationship between holographic scattering and nonminimal surfaces identified in Caminiti et al. (2024). Starting with the observation that the 2-to-2 connected wedge theorem (CWT) admits a converse in pure AdS3, we established a more general connection between holographic scattering and nonminimal surfaces in pure states for any spacetime where the geometry near the boundary is locally AdS3. This extends the $u<d$ result of Caminiti et al. (2024).
We emphasize that our general arguments apply well beyond the static BTZ and conical defect solutions considered in Caminiti et al. (2024). Other nontrivial solutions where our results would apply include, e.g., spinning defects Miskovic and Zanelli (2009); Compère and Fiorucci (2018); Edelstein et al. (2011); Briceño et al. (2021); Li (2024) (see appendix B), spinning BTZ black holes Banados et al. (1992, 1993), and orbiting or colliding defects and black holes Balasubramanian and Ross (2000). Further, in these spacetimes, our scattering setup can be generalized beyond that considered in Caminiti et al. (2024) and illustrated in figure 2, in which $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ have equal size and lie on the same time-slice. Our general arguments allow us to choose the input regions $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ to be any spacelike separated causal diamonds on the boundary cylinder, where the output regions $\hat{\mathcal{R}}_{1}$ and $\hat{\mathcal{R}}_{2}$ are the causal diamonds extending between the future tips of these inputs. Consequently, $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ need not be the same size or lie on a constant time slice, e.g., see figure 4.
@@
Maximal outputs: A key feature of the $u<d$ construction is that the output regions $\hat{\mathcal{R}}_{i}$ are chosen to be maximal, given fixed inputs $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$. We emphasize that if the outputs are smaller, the $u<d$ relation need not imply scattering. In particular, consider sliding the endpoints of $\hat{\mathcal{R}}_{i}$ up along the $\hat{\mathcal{J}}^{-}(r_{i})$ lightcones to produce smaller $\hat{\mathcal{W}}_{i}\subset\hat{\mathcal{R}}_{i}$ – for reference, see figure 4. It turns out that shrinking $\hat{\mathcal{W}}_{i}$ too far may result in an empty scattering region. A dramatic illustration of this situation is given by considering $\hat{\mathcal{V}}_{1}$ and $\hat{\mathcal{V}}_{2}$ to be of equal size, on the boundary of a conical defect geometry with M>4/9M>-4/9. If we set $\hat{\mathcal{W}}_{1}={r_{1}}$ and $\hat{\mathcal{W}}_{2}={r_{2}}$, then holographic scattering is impossible for all choices of $\theta$ and $x$, even when we have $u<d$ Caminiti et al. (2024).
@@
Closed timelike curves: Another notable feature of the $u<d$ result is that in the spinning defect backgrounds, the $u$ surface may probe a region of the geometry with closed timelike curves, even if the scattering itself happens outside of the CTC region. [^31] In the notation of appendix B, this arises when $\theta+x\in(\Delta\phi_{CTC},\pi/b_{+})$ and $\theta<\Delta\phi_{CTC}$. In physical scenarios, one expects that the spinning defect solution describes the spacetime only outside of a localized matter source, such as the spinning string of Maxfield and Wang (2022), and no CTC region actually exists. Hence, in these physical spacetimes, we expect that the RT surfaces will encounter the matter sources in certain regimes, disrupting the $u<d$ connection to holographic scattering, similarly to the situation in section 4. Still it would be interesting to understand, from a bulk reconstruction point of view, what it means for the $u$ entanglement wedge to probe the CTC region, and why this appears to be useful for diagnosing holographic scattering.
@@
Specifically, in geometries which are locally AdS3 outside of certain matter sources, the $u$ surface reaches deeper and deeper into the bulk as $\theta+x$ increases, and it may become disrupted by the presence of bulk matter at a critical value of $\theta+x$ which we call $\Delta\phi_{s}$. For static defect geometries with M<1/4M<-1/4, Caminiti et al. (2024) found $\Delta\phi_{s}=\frac{\pi}{\sqrt{|M|}}$, and for spinning defect geometries with $M<-\frac{1}{4}-J^{2}$, we generalized this to
@@
Our goal in studying the shell geometry in section 4 was to further examine the effects of matter on the $u<d$ converse. In this case, we have $\Delta\phi_{s}=\frac{2}{\mu}\tan^{-1}\frac{\mu}{R}$, and we found examples where holographic scattering is controlled by a naive extension of the $u=d$ formula and yet $\theta+x>\Delta\phi_{s}$. Hence in these situations, the $u$ surface does not exist and so cannot be said to control the scattering. This reinforces the result that bulk matter disrupts the tight relationship between extremal surfaces and bulk scattering introduced in Caminiti et al. (2024).
@@
Conversely, one might conjecture that whenever $u<d$ fails to control scattering, the $u$ surface does not exist at the scattering threshold. This is almost implied by our discussion of eq. (85), except for the possibility that the $\theta+x=\Delta\phi_{s}$ curve intersects the scattering curve not only at eq. (85) but also at some $\theta$ above $\Delta\phi^{*}$, which is the regime where scattering is no longer controlled by the naive extension of the $u=d$ formula (see figure 16). To validate this conjecture in cases where $R$ is very large ($R\gg 1$), or the shell is very light ($\mu\approx 1$), note that $\Delta\phi_{s}\approx\Delta\phi^{*}$ in these limits. [^33] To be more quantitative: for $R>1$, we find $\Delta\phi_{s}-\Delta\phi^{*}<\pi/20\simeq 0.157$ for all choices of $\mu$. The same bound holds with $\mu>0.9$ and any choice of $R$. Hence, for $\theta>\Delta\phi^{*}$, the $u$ surface only exists ($\theta<\Delta\phi_{s}-x$) within a tiny island of parameter space near $x=0$, and it is reasonable to expect that this island does not extend to the scattering threshold curve. Turning now to cases where $R$ is small and the shell is heavy, we observe that in these cases, the (extended) $u=d$ curve provides a lower bound on the scattering threshold, since the shell introduces time delays which make scattering more difficult compared to scattering in the $R=0$ geometry. [^34] Interestingly, this feature does not hold for all parameters; with the introduction of a sufficiently large shell, it is possible that scattering becomes easier – see figure 16. This is because, roughly, light rays travel faster through pure AdS3 than they do in the conical defect geometry, and so a large pure AdS3 region can compensate for the time delays introduced at the shell interface. In these cases, the conjecture is validated by eq. (85) being less than $\Delta\phi^{*}$. To fully validate this conjecture, one must be more careful about intermediate values of $R$, $\mu$, but it appears to hold on a case-by-case basis (e.g., $R=0.75$, $\mu=0.4$).